# Astrophysical Reaction Rates for Charged-Particle Induced Reactions on Proton-Rich Nuclides

### v1

@@ Abstract
Astrophysical reaction rates for reactions with proton-rich nuclides from stability to the proton dripline were calculated with an updated version of the SMARAGD statistical model (Hauser-Feshbach) code. Here, the focus was on reactions with protons or $\alpha$ particles as required for nucleosynthesis in proton-rich matter. For completeness, also neutron-induced reactions are provided for the same set of targets. Some comments on dependencies of rates on various nuclear properties and on the appropriate way to compare to experiments are given. The new rate set for charged-particle induced reactions provides a better description of experimental data than previously widely used rates, especially for reactions involving $\alpha$ particles.
@@ 1 Introduction
Nuclear reaction rates provide the basis for all astrophysical investigations concerning energy generation and nucleosynthesis in the Early Universe as well as in hydrostatic and explosive nuclear burning of stars and stellar objects. They are derived from experimentally measured and/or theoretically predicted nuclear reaction cross sections. A direct experimental determination of reaction cross sections at the interaction energies relevant to astrophysical environments is possible only if the involved nuclei are accessible for reaction studies, i.e., are stable or have sufficiently long half-lives, cross sections are sufficiently large, and thermally excited states of the target nuclei do not contribute considerably to the reaction rate in an astrophysical plasma. Therefore the majority of astrophysical reaction rates used in astrophysical simulations are based on theory and may rely only indirectly on experimentally obtained information required to inform the theoretical treatment of the nuclear properties used in the reaction models. Depending on the relative energy and the level density of the interacting nuclei, different reaction models are employed. For target nuclides of intermediate and high mass number ($A\gtrapprox 20$) the relevant excitation energy of the compound system formed from the combination of projectile and target in most cases is high enough to be located in the regime of unresolved, overlapping resonances. This is the domain of the statistical model of nuclear reactions, originally proposed by Bohr [1] and later refined by [2, 3]. At low level density at the compound formation energy or at high interaction energy, individual resonances and direct reaction mechanisms have to be considered, respectively (see, e.g., [4, 5] for an overview).
@@ 2 Definition of Astrophysical Reaction Rates
The quantity $\sigma^{*}(E,T)$ is the stellar reaction cross section [8, 4, 5], including reactions not only on the ground states of the interacting nuclei but also on their excited states which are populated by thermal excitation in the plasma,
@@
3 The Statistical Model (Hauser-Feshbach)
To assure the applicability of the statistical model of nuclear reactions, the nuclear level density at the compound-nucleus formation energy has to be high enough to permit the use of averaged resonance widths. For the majority of reactions of astrophysical interest across the nuclear chart, it has been shown that the statistical model is applicable to the calculation of astrophysical reaction rates [17]. Exceptions are target nuclides with low mass number, at shell closures (for rates at low $T$), and towards the driplines. A large number of resonances at the compound formation energy implies that the sum in Equation (3) extends into the unresolved resonance region. Therefore above the excitation energy $E^{\mathrm{x}}_{\mathrm{last}}$ of a last resolved level, an integration over the density $\rho$ of nuclear levels with spin $J$ and parity $\Pi$ has to be invoked:
@@ 4 Implementation in SMARAGD
Masses were taken from the 2020 Atomic Mass Evaluation (AME2020) [22, 23]. Experimentally unknown masses (and microscopic corrections for the calculation of the nuclear level density) were taken from [24]. Experimentally known spins and parities of ground states and low-lying states were taken from NuDat 2.8 [25]. Up to 40 levels were included but only up to the last level for which an unambiguous assignment of spin and parity is present and up to which the level scheme can be considered as being complete. Above the last included discrete level with energy $E_{\mathrm{last}}^{\mathrm{x}}$, a theoretical nuclear level density description was employed. The description is based on the approach described in [17, 11] combining a constant temperature formula for low excitation energies with a backshifted Fermi-gas formula having a level density parameter that depends on the excitation energy [17]. In this previous description, the parities were assumed to be equi-distributed. Here, a parity distribution also depending on the excitation energy is implemented [26].
@@ 6.1 Some comments on comparing data to predictions
The use of experimental data is always preferable to using theoretical predictions and a reaction rate set for implementation in nuclear network calculations usually includes a mix of experimental and theoretical reaction rates. However, it is important to realise that only a small subset of the required rates, especially for nuclides in the intermediate and heavy mass range, may be fully constrained by data. On one hand, reaction cross-sections below the Coulomb barrier become tiny and are difficult or currently impossible to measure, even for stable targets. On the other hand, a measurement of g.s. cross sections, as usually performed in laboratory experiments, may only constrain a small contribution to the astrophysical reaction rate because the contributions of reactions on excited states of the target nuclei often dominate the rate, especially at the elevated temperatures encountered in explosive nucleosynthesis [4, 19, 43, 44]. In such cases, the results of the measurement have to be supplemented or replaced by theory, with the data only providing a guideline for assessing the validity of the theoretical treatment of the g.s. contribution. The online tables also include the g.s. contribution to the total astrophysical rate as guidance.
@@
Last but not less important is the consideration of the applicability of the model. The statistical model assumes that there is a ”sufficiently large” number of overlapping resonances in the compound nucleus at the compound formation energy. When the nuclear level density at the compound formation is too low, the calculated model cross section will overestimate the actual cross section because inherent in the model is a summation over an extended number of spins assumed to be present at the relevant excitation energy of the compound nucleus. The inherent nuclear level density is low at shell closure and towards the dripline, the latter because of the lower proton and $\alpha$ separation energies. This means that for lower temperatures (meaning lower compound formation energies) the Hauser-Feshbach model may not be applicable to calculate reaction rates. This has been investigated previously [17]. In that previous publication, temperature limits for the application have been derived. These limits have been shown in the tables of [11] and are also included in the present tables of reaction rates and their fits. Using values outside their range of applicability it should not be surprising if deviations to, e.g., measured cross sections and rates appear.
@@ 6.2 Comparison to selected experimental data
As explained above, even a good reproduction of experimental cross sections can only increase the confidence in the predictions but does not necessarily imply also accurate reaction rates. This is especially the case when the contribution of thermally excited target states is not negligible. For measurements above the relevant energy window (Gamow window) it has to be assessed carefully whether the measured cross sections and the astrophysical reaction rate have a similar sensitivity to the contributing nuclear properties. At higher interaction energies, the sensitivity may be quite different from the one at astrophysical, lower energies (see [42, 4] for examples).
@@ 7 Comparison to calculations with different treatment of nuclear properties
Figure 13 shows the impact of the different choice of optical proton+nucleus potential. It is evident that the impact is much smaller than the one of the changed $\alpha$+nucleus potential but nevertheless the modification was necessary to obtain better agreement with experimental reaction cross sections.

### v2

@@ Abstract
Astrophysical reaction rates for reactions with proton-rich isotopes of Ne to Bi from stability to the proton dripline were calculated with an updated version of the SMARAGD statistical model (Hauser-Feshbach) code. Here, the focus was on reactions with protons or $\alpha$ particles as required for nucleosynthesis in proton-rich matter. For completeness, also neutron-induced reactions are provided for the same set of targets. Some comments on dependencies of rates on various nuclear properties and on the appropriate way to compare to experiments are given. The new rate set for charged-particle induced reactions provides a better description of experimental data than previously widely used rates, especially for reactions involving $\alpha$ particles.
@@ 1 Introduction
Nuclear reaction rates provide the basis for all astrophysical investigations concerning energy generation and nucleosynthesis in the Early Universe as well as in hydrostatic and explosive nuclear burning of stars and stellar objects. They are derived from experimentally measured and/or theoretically predicted nuclear reaction cross sections. A direct experimental determination of reaction cross sections at the interaction energies relevant to astrophysical environments is possible only if the involved nuclei are accessible for reaction studies, i.e., are stable or have sufficiently long half-lives, cross sections are sufficiently large, and thermally excited states of the target nuclei do not contribute considerably to the reaction rate in an astrophysical plasma. Therefore the majority of astrophysical reaction rates used in astrophysical simulations are based on theory and may rely only indirectly on experimentally obtained information required to inform the theoretical treatment of the nuclear properties used in the reaction models. Depending on the relative energy and the level density of the interacting nuclei, different reaction models are employed. For target nuclides of intermediate and high mass number ($A\gtrapprox 20$) the relevant excitation energy of the compound system formed from the combination of projectile and target in most cases is high enough (see Section 3) to apply the statistical model of nuclear reactions, originally proposed by Bohr [1] and later refined by [2, 3]. At low level density at the compound formation energy or at high interaction energy, individual resonances and direct reaction mechanisms have to be considered, respectively (see, e.g., [4, 5] for an overview).
@@ 2 Definition of Astrophysical Reaction Rates
The quantity $\sigma^{*}(E,T)$ is the stellar reaction cross section [16, 8, 4, 5], including reactions not only on the ground states of the interacting nuclei but also on their excited states which are populated by thermal excitation in the plasma,
@@
In the stellar cross section, the contributions of excited states are weighted by $(2J_{i}+1)(E-E^{\mathrm{x}}_{i})/E=(2J_{i}+1)(1-E^{\mathrm{x}}_{i}/E)$. This looks differently compared to the usual Boltzmann population factors exhibiting an exponential suppression depending on excitation energy $(2J_{i}+1)\exp(-E^{\mathrm{x}}_{i}/(k_{\mathrm{B}}T))$. The stellar cross section in this form is obtained when transforming the Boltzmann-weighted sum of reaction rate integrals for each excited state to the single integral shown in Equation (LABEL:eq:rate). This not only involves an exchange of summation and integration but also requires a substitution of the integration variables, leading to the transformed weights [16, 8, 4, 5]. The use of the stellar cross section not only simplifies the expression for the reactivity but also allows to judge the importance of excited state contributions better than by just examining Boltzmann population factors multiplied by integrals. The single integration of a stellar cross section further provides better numerical accuracy in the calculation of a rate and that is why it has been used in theoretical approaches since [16] (see, e.g., [8, 9, 10, 11, 13, 4]).
@@
Depending on the nucleus and the temperature, the sum in Equation (3) has to be extended into a region of excitation energy with insufficient knowledge of discrete excited states. In this case, above the excitation energy $E^{\mathrm{x}}_{\mathrm{last}}$ of a last discrete level included, an integration over the density $\rho$ of nuclear levels with spin $J$ and parity $\Pi$ has to be invoked:
@@
3 The Statistical Model (Hauser-Feshbach)
To assure the applicability of the statistical model of nuclear reactions, the nuclear level density at the compound-nucleus formation energy has to be high enough to assume overlapping resonances that permit the use of averaged resonance widths. As can be seen from the sum appearing in Equation (5) below, an additional assumption is that in principle states with all spins and parities are present at the compound formation energy. The statistical model cross section is smooth without exhibiting resonance features. Due to the fact that the calculation of the reaction rate involves an integration over an energy range, i.e., an energy average of cross sections (see Equation LABEL:eq:rate), the criterion for the applicability of the statistical model is more relaxed when applied to stellar rates than to reaction cross sections. A smaller number of contributing resonances can be sufficient when applying the energy average of a statistical model cross section because resonance features in the actual cross section are smoothed out, anyway [17]. For the majority of reactions of astrophysical interest across the nuclear chart, it has been shown that the statistical model is applicable to the calculation of astrophysical reaction rates [17]. Exceptions are target nuclides with low mass number, at shell closures (for rates at low $T$), and towards the driplines.
@@ 4 Implementation in SMARAGD
Masses were taken from the 2020 Atomic Mass Evaluation (AME2020) [22, 23]. Experimentally unknown masses (and microscopic corrections for the calculation of the nuclear level density) were taken from [24]. Experimentally known spins and parities of ground states and low-lying states were taken from the 2021 ENSDF+Nubase snapshot as provided by RIPL-3 [37]. Up to 40 levels were included but only up to the last level for which an unambiguous assignment of spin and parity is present and up to which the level scheme can be considered as being complete. Above the last included discrete level with energy $E_{\mathrm{last}}^{\mathrm{x}}$, a theoretical nuclear level density description was employed. The description is based on the approach described in [17, 11] combining a constant temperature formula for low excitation energies with a backshifted Fermi-gas formula having a level density parameter that depends on the excitation energy [17]. In this previous description, the parities were assumed to be equi-distributed. Here, a parity distribution also depending on the excitation energy is implemented [26].
@@ 6.1 Some comments on comparing data to predictions
The use of experimental data is always preferable to using theoretical predictions and a reaction rate set for implementation in nuclear network calculations usually includes a mix of experimental and theoretical reaction rates. However, it is important to realise that only a small subset of the required rates, especially for nuclides in the intermediate and heavy mass range, may be fully constrained by data. On one hand, reaction cross-sections below the Coulomb barrier become tiny and are difficult or currently impossible to measure, even for stable targets. On the other hand, a measurement of g.s. cross sections, as usually performed in laboratory experiments, may only constrain a small contribution to the astrophysical reaction rate because the contributions of reactions on excited states of the target nuclei often dominate the rate, especially at the elevated temperatures encountered in explosive nucleosynthesis [4, 19, 43, 44]. In such cases, the results of the measurement have to be supplemented or replaced by theory, with the data providing the means for assessing the validity of the theoretical treatment only of the g.s. contribution. The online tables also include the g.s. contribution to the total astrophysical rate as guidance.
@@
The fact that the widths depend on the relative energy (and also the spin of the involved levels) has the important consequence that the excited state contributions to the stellar rate may exhibit different sensitivities than the contribution from reactions on the g.s. of a nucleus. This, in turn, implies that a comparison between data and theory for g.s. reactions may not provide the required information to improve the prediction of the stellar rate. This is especially relevant at the elevated temperatures of explosive nucleosynthesis with their tiny g.s. contributions to the stellar rate (see the online tables and Refs. [4, 19, 43, 44]).
@@
Last but not less important is the consideration of the applicability of the model. The statistical model assumes that there is a ”sufficiently large” number of overlapping resonances in the compound nucleus at the compound formation energy. When the nuclear level density at the compound formation is too low, the calculated model cross section will overestimate the actual cross section because inherent in the model is a summation over an extended number of spins assumed to be present at the relevant excitation energy of the compound nucleus. The inherent nuclear level density is low at shell closure and towards the dripline, the latter because of the lower proton and $\alpha$ separation energies. This means that for lower temperatures (meaning lower compound formation energies) the Hauser-Feshbach model may not be applicable to calculate reaction rates. This has been investigated previously [17]. In that previous publication, temperature limits for the application have been derived. These limits have been shown in the tables of [11] and are also included in the present tables of reaction rates and their fits. Using values outside their range of applicability it is to be expected that deviations to, e.g., measured cross sections and rates, increase.
@@ 6.2 Comparison to selected experimental data
As explained above in Section 6.1, even a good reproduction of experimental g.s. cross sections can only increase the confidence in the predictions but does not necessarily imply also accurate reaction rates. This is especially the case when the contribution of thermally excited target states is not negligible because, as explained above, reactions on excited target states may be sensitive to other widths than reactions on the ground state. Furthermore, for measurements above the relevant energy window (Gamow window) it has to be assessed carefully whether the measured cross sections and the astrophysical reaction rate have a similar sensitivity to the contributing nuclear properties. At higher interaction energies, the sensitivity may be quite different from the one at astrophysical, lower energies (see [42, 4] for examples).
@@ 7 Comparison to calculations with different treatment of nuclear properties
Figure 14: Comparison of rates obtained with and without the inclusion of experimental excited levels for $\alpha$-induced reactions.
Figure 15: Comparison of rates obtained with and without the inclusion of experimental excited levels for proton-induced reactions.
@@
Figure 13 shows the impact of different choices of the optical proton+nucleus potential. It is evident that the impact is much smaller than the one of the changed $\alpha$+nucleus potential but nevertheless the modification was necessary to obtain better agreement with experimental reaction cross sections [29, 30, 31] (see also Section 4). As shown in Figure 8, the modified potential also fares well in comparison to the data of Refs. [48, 49, 50] which were not available when the modification was suggested originally.